Ticket Number: HEA-NOC/20020820-01 Ticket Status: UPDATE
Ticket Type: Unscheduled Ticket Source: HEA-NOC
Ticket Scope: Site Site/Line: Uther
Ticket Owner: [log in to unmask] Problem Fixer: Lancomms
Ticket Opened: Problem Start: 20020820 07:30 UTC+1
Ticket Closed: Problem End:
Tracebacks in the log on Dublin ATM access router Uther.hea.net.
Traffic to ATM connected institutions at the Dublin PoP (TCD, UCD,
ITnet, DCU, DIT, NUI Maynooth, DIAS, HEA, NCAD) diverted to backup
ATM router for a few minutes at 07:30 and 08:30, with impeded
Reported to LAN Comms, ticket #lc00-5d7bpe, who have in-band access
and are investigating. Hardware fault suspected at this time.
20020820 10:00 davew
Further investigation by LAN Comms suggests a software problem.
Case is being opened with Cisco, have asked them to raise priority
as we are during the CAO critical period.
20020820 12:20 davew
Another failure at 12:09. Cisco believe it is a hardware fault and
are supplying are replacement ATM card for the router; swapout will
be scheduled shortly.
20020820 15:06, 15:19, 15:31 AH
Repeated failures. Customers reroute via backups during failures.
20020820 17:30 davew
Further repeated failures, some lengthy. Shutting down ATM interface
and failing over to backup router. There should be sufficient capacity
on this router for overnight traffic; NOC will continue to monitor and
intervene again if necessary in advance of tomorrow's card swap
20020827 gs Lancomms on site, slave rsp removed, suspected hardware fault. Monitoring overnight. Traffic continues to run at reduced capacity on the backup.
20020829 mn RSP card replaced as advised by Cisco/LAN Comms. Traffic will be re-routed via Uther.
20020829 08:30 gs Traffic re-routing. Ticket will remain open for monitoring.
Time to Fix:
Not yet known.